A new reserve is an equilateral triangle with sides 200 m. It is to be surfaced with instant turf. Find the area of turf needed to grass it.
Answer by checkley71(8403)

AREA-1/2BH THE BASE=200M
THE HEIGHT IS FOUND USING THE PYTHAGOREAN THEORUM A^2+B^2=C^2
A=200 & B=100
200^2+100^2=C^2
40000+10000=C^2
50000=C^2
C=SQRT50000
C=223.61M. THIS IS THE HEIGHT (H).
AREA=.5*100*223.61
AREA=11,180.5 M^2